OpenStudy (anonymous):

Use the midpoint formula repeatedly to find the three points that divide the segment joining (x₁,y₁) and (x₂,y₂) into four equal parts. What does this mean?

OpenStudy (anonymous):

cut it in half, then cut it in half again

OpenStudy (anonymous):

midpoint is \[(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\]

OpenStudy (anonymous):

now find the midpoint of \[(x_1,y_1)\] and \[(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\]

OpenStudy (anonymous):

so just leave the symbols in it?

OpenStudy (anonymous):

that should be fun. yes if they didn't give you numbers what else can you do right?

OpenStudy (anonymous):

yeah, okay. thanks